# Curvature-Constrained Vector Field for Motion Planning of Nonholonomic Robots

### v1

@@ IV Saturated Control Laws with Dynamic Gain
In [32], the angular velocity for the Dubins-car-like robots to track the VF is also subject to saturation. However, the orientation error is shown to be stabilized only when the angular velocity does not saturate persistently. To achieve this, implicit constraints on the system dynamics are introduced. However, it remains unclear how to establish these constraints in practice. In contrast, our approach offers a straightforward criterion for the co-design of the CVF and control laws. Under the proposed control inputs, the nonholonomic robot exits the saturation region under curvature constraint and hence the angular velocity would not saturate persistently, leading to the results that the orientation error is monotonically stabilized.
According to (9), the first term on the left-hand side of (25) reflects the distance between the singular point $\bm{p}_{\delta}$ and the region $\mathcal{A}_{i=2,3}$, where the flow fields are blended. The second term depicts the range of the blending region $\mathcal{A}_{i=2,3}$. To ensure the existence of dynamic gain which minimizes the saturation region to $\mathcal{S}$, the changing rate in CVF’s orientation should be limited by the saturation bound $\bar{\omega}$ outside $\mathcal{S}$. When the blending region is close to the singular point and hence the CVF changes rapidly and the first term is large, the range of the blending region should be enlarged to reduce the effect of singularity, which requires a smaller second term.
By co-design of the CVF and control laws, the orientation error is stabilized in Theorem 4. In this way, the nonholonomic robot tracks the CVF without violating the curvature constraint. Based on these results, we further employ the input-to-state stability (ISS) to show the almost global convergence of the robot’s configuration in the following theorem.
@@
In the literature, most VF-based motion planning algorithms for nonholonomic robots with curvature constraints focus on incorporating the curvature constraint either in the design of the VF [34, 36, 37] or in the derivation of control laws [32]. Consequently, it remains unclear whether these robots can reliably follow the VFs under limited actuation. We fill this gap in the paper by integrating the curvature constraint into the co-design of the CVF and saturated control laws with dynamic gain, providing rigorous mathematical guarantee on the convergence of the nonholonomic robot under curvature constraint.
@@
Figure 6: Simulation results validate Theorem 5. Denote the simulation time for the $i$-th example as $t_{i}$ and the terminal time as $t_{i,f}$. The subfigures LABEL:sub@fig:Thm4_Set_1_errorLABEL:sub@fig:Thm4_Set_2_error show that the configuration in each example converges to the target configuration $\bm{g}_{d}=(x_{d},y_{d},\theta_{d})$ under the control inputs. The subfigures LABEL:sub@fig:Thm4_Set_1_t0-LABEL:sub@fig:Thm4_Set_1_t3 present snapshots of Example 1-3 at different time instances. The subfigures LABEL:sub@fig:Thm4_Set_2_t0-LABEL:sub@fig:Thm4_Set_2_t3 are snapshots of Example 4-7 at different time instances.
@@ V-A2 Results
We begin with the saturation behavior depicted in Theorem 3. In Figures 4(a) and 4(b), the unsaturated angular velocity $\omega_{0}$ in (20) is shown by gray dotted line. It exceeds the bound $\bar{\omega}$ presented by red dotted line and thereby the angular velocity saturates when the robot is in the saturation region $\mathcal{S}=\left\{\bm{g}\mid r_{\delta}<\rho\right\}$. While in Figures 4(c)-LABEL:sub@fig:Lemma2_Exp7, there always holds that $\omega=\omega_{0}$ and hence the angular velocity does not saturate. The above phenomenon confirms that the angular velocity with dynamic gain only saturates in $\mathcal{S}$. Moreover, the dynamic gain also enables the robot to leave the saturation region in finite time such that the angular velocity would not saturation persistently, as shown by Figures 4(a)-4(c).
Following the above discussion, we illustrate the evolution of the orientation error $\theta_{e}$ in Figure 5. By co-design of the CVF and control laws, we provide the explicit condition (25) in Theorem 4, under which the orientation error $\theta_{e}$ asymptotically converges to zero in all the examples. Therefore, the trajectories of the closed-loop system are guaranteed to converge to the integral curves of the CVF. It should be noted that the initial orientation of the robot in Example 7 aligns with the CVF, so that the orientation error $\theta_{e}$ remains zero.
@@ V-B1 Setup
In this subsection, we conduct 1000 trials of Monte Carlo experiments with different initial and target configurations to compare the performance of the proposed CVF method, with respect to other VF-based motion planning algorithms with feedback control laws for nonholonomic robots. The benchmarks include the dynamic vector field (DVF) [35, Eq.20, Eq.24], the dipole-like attractive vector field (AVF) [29, Eq.8, Eq.18], the curvature-constrained Lyapunov vector field (CLVF) [36, Eq.16 Eq.50], and the guiding vector field (GVF) [31, Eq.4, Eq.24]. Among these candidates above, the DVF and AVF guide the unicycle to the target configuration, while CLVF and GVF navigate the UAV to follow a target circular path, where both scenarios are within the scope of the proposed approach. The comparisons are divided into two categories, including comparisons to DVF and AVF on the unicycle and comparisons to GVF and CLVF on UAV.
For the unicycle with linear velocity bounded by $v_{\text{max}}=3$ and $v_{\text{min}}=0$, we prepare four sets of target configurations, where the target positions are evenly distributed on the circle $x^{2}+y^{2}=r_{2}^{2}$ and the target orientation is tangent to the circle. For UAV, the linear velocity is chosen to be constant at $v_{\text{min}}=3$ and the target positive limit set is specified by $\mathcal{L}_{d}^{+}=\left\{\|\bm{p}\|=r_{2},\theta=\varphi+\pi/2\right\}$, where four sets of target configurations are also evenly distributed. The maximum curvature is set as $\bar{\kappa}=1$, the minimal turning radius $\rho=1$, and the parameters of the proposed method are the same as in the previous subsection. We conduct 250 trials of Monte Carlo simulation for each target configuration, 1000 trials for each candidate in total, with initial configuration randomly distributed on $x_{0}\in[-15\rho,15\rho]$, $y_{0}\in[-15\rho,15\rho]$ and $\theta_{0}\in[0,2\pi)$.
@@ V-B2 Metrics
The metrics for comparisons are divided into two categories to show the quality of the reference trajectory and the control performance, respectively. To evaluate the quality of the reference trajectories, we compute the percentage of trials where the integral curves have curvature bounded by $\bar{\kappa}$. In addition, the relative length of the integral curves is defined as $L_{r}=L/\|\bm{p}_{d}-\bm{p}_{0}\|$, where $L$ is the length of the integral curve from initial position $\bm{p}_{0}$ to the target position $\bm{p}_{d}$. As for the control performance, the percentage of trials where the control inputs satisfy the curvature constraint (2) is counted. Next, the average curvature of the trajectories over time is computed to reflect the feasibility of the control inputs w.r.t. the capability of the nonholonomic robot. In addition, we record the average time consumption for the robot to converge to the target configuration. Finally, we define the root-mean-square error (RMSE) between $\{\omega(t_{i})\}$ and $\{\omega(t_{i+1})\}$ to measure the acceleration of control inputs, where $t_{i}$ is the $i$-th time steps in the simulation.
@@
Metrics CVF (Unicycle) DVF [35] AVF [29] CVF (UAV) CLVF [36] GVF [31]
Trajectory Quality Percentage of $\kappa\leq\bar{\kappa}$ 1.$0$ 0.9960 1.0000 1.0000 0.8050
@@ V-B3 Results
The comparison results are presented in Table II. It can be seen that the proposed method outperforms the other algorithms by providing control inputs that satisfy the curvature constraint for all trials.
For the results of the unicycle, the DVF is not taken into account for the comparison of trajectory quality since it is a map from $\mathbb{R}^{2}\times\mathbb{S}$ to $\mathbb{R}^{2}$ and hence the integral curves are not the reference trajectories that we consider in this paper. Throughout 1000 trials, the control inputs provided by DVF and AVF contain sharp turns that exceed the capability of the unicycle since the curvature constraint is not addressed in [35, 29], while the proposed CVF-based algorithm provides reference trajectories with control inputs that both satisfy the curvature constraint under the guarantees claimed in Theorems 2 and 5.
@@ VI Hardware Experiments
As mentioned in Section II, the proposed method under the problem formulation is suitable to a wide range of nonholonomic robots. In the real-world implementations presented in this section, we prepare two distinct robotic systems, namely the Ackermann UGV and the fixed-wing UAV, to further illustrate the effectiveness of the proposed method.
@@ VI-A1 Setup
[…]
Similar to the setup in the previous section, we prepare two sets of independently conducted examples for the Ackermann UGV. In the first set of three examples, the UGV starts in or passes through the saturation region $\mathcal{S}$, while the UGV stays outside $\mathcal{S}$ for the four examples in the second set. The initial and target configurations are summarized in Table III. In order to generate control inputs that start from zero, we specify the linear velocity gain by $k_{v}=(1-e^{-0.3t})v_{\text{max}}$, where $t$ is the run time and other control parameters are the same as those in the previous section.
@@ VI-A2 Results
The experiment results are presented in Figure 8. Figures 8(a)LABEL:sub@fig:Exp_inputs4567 shows the tracking results of the control inputs. The errors between the current and target configurations are plotted in Figure 8(c). The snapshots of Examples 1-3 and Examples 4-7 are presented in Figure 8(d)LABEL:sub@fig:Exp_half_123LABEL:sub@fig:Exp_full_123 and Figure 8(e)LABEL:sub@fig:Exp_half_4567LABEL:sub@fig:Exp_full_4567, respectively. Due to the measurement noises, we observe that the tracking results of the linear velocity contain small fluctuations around the desired value. In angular velocity, the track error becomes larger due to the parameter uncertainties and response time of the UGV’s chassis. However, despite the real-world disturbances above, the Ackermann UGV consistently converges to the target configuration with bounded trajectory curvature in all examples. The experimental results showcase the effectiveness and robustness of the feedback motion planning algorithm in real-world applications.
(a) Control inputs tracking results of Exp 1-3.
@@
Figure 8: Experimental results with the Ackermann UGV. The tracking errors of the control inputs are presented by the shaded areas around the desired values in subfigures LABEL:sub@fig:Exp_inputs_123LABEL:sub@fig:Exp_inputs4567.
@@ VI-B HIL Experiments with the Fixed-Wing UAV
Figure 9: HIL experiment platform same as [40] except the planning algorithm.
@@ VI-B1 Setup
To demonstrate that the proposed algorithm given by closed-form expressions can be implemented in real-time onboard computation, we conduct the hardware-in-the-loop (HIL) experiments with the fixed-wing UAV. The hardware-in-the-loop (HIL) system is illustrated by Figure 9. In this architecture, the onboard computer handles all the computation of motion planning algorithm, the autopilot serves as the flight control unit and generates commands for actuators based on our planned control inputs, and the simulator mirrors the real-world flight environment and provides configuration feedback.
[…]
The HIL experiments are conducted in the following procedure. At the beginning, the fixed-wing UAV takes off and flies to the initial configuration $\bm{g}_{0}$, and then the motion planning algorithm is activated, guiding the UAV to the target configuration $\bm{g}_{d}$. After the UAV arrives at $\bm{g}_{d}$, the motion planning algorithm is deactivated until the autopilot drives the UAV to the next initial configuration without landing. This process is repeated for all the examples in Table IV, where all the target configurations are contained by a common target positive limit set $\mathcal{L}_{d}^{+}=\{\bm{g}\mid r=r_{2},\ \theta=\theta_{r}\}$.
@@ VI-B2 Setpoint Derivation
The control inputs of linear velocity and angular velocity in 2D provided by our algorithm cannot be directly applied to 3D fixed-wing UAVs. Therefore, we first convert the linear and angular velocities into the setpoints of attitude and thrust for the autopilot platform as below.
For simplicity, we assume that the attack angle is sufficiently small and there is no slide-slip angle, and hence the body-fixed frame and velocity frame of the fixed-wing UAV are equivalent. The setpoint for the UAV attitude is given by the yaw-pitch-roll Euler angles, denoted by ($\alpha$, $\beta$, $\gamma$). The attitude matrix $\bm{R}\in\mathrm{SO}(3)$ is given by
@@
where $K_{\beta}$ is the gain, $h$ is the current flight height and $V$ is the airspeed. When the experiment begins, the UAV takes off and then circles in the horizontal plane where $h=h_{d}$, waiting for the motion planning algorithm to take over. In this case, we assume that the pitch angle $\beta$ is already stabilized to zero in the following derivation.
@@
Subsequently, we aim to track the planned angular velocity $\omega$ with the time derivative of the yaw angle. In accordance with the coordinated turn constraint
$\dot{\alpha}=-\frac{g}{V}\tan{\gamma}\cos{\beta}\overset{\mathrm{\beta=0}}{=}-\frac{g}{V}\tan{\gamma},$
we derive the setpoint of roll angle for $\dot{\alpha}$ to follow $\omega$ as
$\gamma_{d}=-\arctan\frac{\omega V}{g},$ (29)
where $g$ is the gravitational acceleration and $\omega$ is specified by (22). Finally, to track the planned linear velocity $v_{x}$ given by (17) with airspeed $V$, we assign the thrust by

### v2

@@ IV Saturated Control Laws with Dynamic Gain
In [yao2023guiding], the angular velocity for Dubins-car-like robots to track the VF is also subject to saturation. However, the orientation error is shown to be stabilized only when the angular velocity does not saturate persistently. To achieve this, implicit constraints on the system dynamics are introduced. However, it remains unclear how to establish these constraints in practice. In contrast, our approach offers a straightforward criterion for the co-design of the CVF and control laws. Under the proposed control inputs, the nonholonomic robot exits the saturation region under curvature constraint, and hence the angular velocity would not saturate persistently, leading to the results that the orientation error is monotonically stabilized.
According to the decomposition of workspace in (9), the first term on the left-hand side of (25) reflects the distance between the singular point $\bm{p}_{\delta}$ and the region $\mathcal{A}_{i=2,3}$, where the flow fields are blended. The second term depicts the range of the blending region $\mathcal{A}_{i=2,3}$. To ensure the existence of dynamic gain that minimizes the saturation region to $\mathcal{S}$, the changing rate in CVF’s orientation should be limited by the saturation bound $\bar{\omega}$ outside $\mathcal{S}$. When the blending region is close to the singular point and hence the CVF changes rapidly and the first term is large, the range of the blending region should be enlarged to reduce the effect of singularity, which requires a smaller second term.
Through the co-design of the CVF and control laws, the orientation error is stabilized as stated in Theorem 4. In this way, the nonholonomic robot tracks the CVF without violating the curvature constraint. Building on these results, we further utilize input-to-state stability (ISS) to demonstrate the almost global convergence of the robot’s configuration in the following theorem.
@@
In the literature, most VF-based motion planning algorithms for nonholonomic robots with curvature constraints focus on incorporating the curvature constraint either in the design of the VF [lau2015fluid, pothen2017curvature, shivam2021curvature] or in the derivation of control laws [yao2023guiding]. Consequently, it remains unclear whether these robots can reliably follow the VFs under limited actuation. We fill this gap in the paper by integrating the curvature constraint into the co-design of the CVF and saturated control laws with dynamic gain, providing rigorous guarantee on the convergence of the nonholonomic robot under curvature constraint.
@@
Figure 7: Simulation results validate Theorem 5. Denote the simulation time for the $i$-th example as $t_{i}$ and the terminal time as $t_{i,f}$. The subfigures LABEL:sub@fig:Thm4_Set_1_error, LABEL:sub@fig:Thm4_Set_2_error show that the configuration in each example converges to the target configuration $\bm{g}_{d}=(x_{d},y_{d},\theta_{d})$ under the control inputs. The subfigures LABEL:sub@fig:Thm4_Set_1_t0-LABEL:sub@fig:Thm4_Set_1_t3 present snapshots of Example 1-3 at different time instances. The subfigures LABEL:sub@fig:Thm4_Set_2_t0-LABEL:sub@fig:Thm4_Set_2_t3 are snapshots of Example 4-7 at different time instances.
@@ V-A2 Results
We begin with the saturation behavior depicted in Theorem 3. In Figures 5(a) and 5(b), the unsaturated angular velocity $\omega_{0}$ in (20) is represented by the gray dotted line. It exceeds the bound $\bar{\omega}$ presented by the red dotted line, thereby the angular velocity saturates when the robot is in the saturation region $\mathcal{S}=\left\{\bm{g}\mid r_{\delta}<\rho\right\}$. In contrast, in Figures 5(c)-LABEL:sub@fig:Lemma2_Exp7, the condition $\omega=\omega_{0}$ always holds, implying that the angular velocity does not saturate. Moreover, the dynamic gain also enables the robot to leave the saturation region in finite time such that the angular velocity would not saturation persistently, as shown by Figures 5(a)-5(c).
Following the above discussion, we illustrate the evolution of the orientation error $\theta_{e}$ in Figure 6. By co-designing of the CVF and control laws, the orientation error $\theta_{e}$ asymptotically converges to zero in all examples as stated in Theorem 4. It should be noted that the initial orientation of the robot in Example 7 aligns with the CVF, such that the orientation error $\theta_{e}$ remains zero.
@@ V-B1 Setup
In this subsection, we conduct 1000 trials of Monte Carlo experiments with different initial and target configurations to compare the performance of the proposed CVF method, with respect to other VF-based motion planning algorithms with feedback control laws for nonholonomic robots. The benchmarks include the dynamic vector field (DVF) [he2024simultaneous, Eq.20, Eq.24], the dipole-like attractive vector field (AVF) [Panagou2017Distributed, Eq.8, Eq.18], the curvature-constrained Lyapunov vector field (CLVF) [pothen2017curvature, Eq.16 Eq.50], and the guiding vector field (GVF) [yao2021singularity, Eq.4, Eq.24]. Among these candidates, the DVF and AVF guide the unicycle to the target configuration, while the CLVF and GVF enable the UAV to follow a target circular path, with both scenarios falling within the scope of the proposed approach. The comparisons are categorized into two groups: comparisons between the DVF and AVF on the unicycle, and comparisons between the GVF and CLVF on the UAV.
For the unicycle with linear velocity bounded by $v_{\text{max}}=3$ and $v_{\text{min}}=0$, we prepare four sets of target configurations, where the target positions are uniformly distributed on the circle $x^{2}+y^{2}=r_{2}^{2}$ and the target orientation is tangent to the circle. For UAV, the linear velocity is selected to be constant at $v_{\text{min}}=3$ and the target positive limit set is specified by $\mathcal{L}_{d}^{+}=\left\{\|\bm{p}\|=r_{2},\theta=\varphi+\pi/2\right\}$, where four sets of target configurations are also uniformly distributed. The maximum curvature is defined as $\bar{\kappa}=1$, the minimal turning radius as $\rho=1$, and the parameters of the proposed method remain consistent with those in the previous subsection. We conduct 250 trials of Monte Carlo simulation for each target configuration, totaling 1000 trials for each candidate, with the initial configuration randomly distributed over $x_{0}\in[-15\rho,15\rho]$, $y_{0}\in[-15\rho,15\rho]$, and $\theta_{0}\in[0,2\pi)$.
@@ V-B2 Metrics
The metrics are divided into two categories to compare the reference trajectory quality and the control performance, respectively. To evaluate the quality of the reference trajectories, we compute the percentage of trials in which the integral curves have curvature bounded by $\bar{\kappa}$. In addition, the relative length of the integral curves is defined as $L_{r}=L/\|\bm{p}_{d}-\bm{p}_{0}\|$, where $L$ is the length of the integral curve from initial position $\bm{p}_{0}$ to the target position $\bm{p}_{d}$. Regarding the control performance, the percentage of trials where the control inputs satisfy the curvature constraint (2). Next, the average curvature of the trajectories over time is computed to reflect the feasibility of the control inputs w.r.t. the capability of the nonholonomic robot. Moreover, we record the average time required for the robot to converge to the target configuration. Finally, we define the root-mean-square error (RMSE) between $\{\omega(t_{i})\}$ and $\{\omega(t_{i+1})\}$ to measure the acceleration of control inputs, where $t_{i}$ is the $i$-th time step in the simulation.
@@
Metrics CVF (Unicycle) DVF [he2024simultaneous] AVF [Panagou2017Distributed] CVF (UAV) CLVF [pothen2017curvature] GVF [yao2021singularity]
Trajectory Quality Percentage of $\kappa\leq\bar{\kappa}$ 1.0000 \ 0.9960 1.0000 1.0000 0.8050
@@ V-B3 Results
The comparison results are presented in Table II. It can be observed that the proposed method outperforms the other algorithms by generating control inputs that adhere to the curvature constraint across all trials.
For the results of the unicycle, the DVF is not taken into account for the comparison of trajectory quality since it is a map from $\mathbb{R}^{2}\times\mathbb{S}$ to $\mathbb{R}^{2}$ and hence the integral curves do not represent the reference trajectories we consider in this paper. Throughout 1000 trials, the control inputs provided by DVF and AVF contain sharp turns that exceed the capability of the unicycle since the curvature constraint is not addressed in [he2024simultaneous, Panagou2017Distributed]. In contrast, the proposed CVF-based algorithm provides reference trajectories with control inputs that both satisfy the curvature constraint under the guarantees established in Theorems 2 and 5.
@@ VI Hardware Experiments
As mentioned in Section II, the proposed method under the problem formulation is applicable to a wide range of nonholonomic robots. In the real-world implementations presented in this section, we utilize two distinct robotic systems, namely the Ackermann UGV and the fixed-wing UAV, to further demonstrate the effectiveness of the proposed method.
@@ VI-A1 Setup
To verify the efficacy of the closed-loop motion planning algorithm in real-world applications, we select the Ackermann UGV  [^1] https://github.com/agilexrobotics/limo-doc with minimum turning radius $\rho=$0.6\text{\,}\mathrm{m}$\$ and maximum linear velocity $v_{\text{max}}=$1\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ to conduct experiments. The experimental setup is illustrated in Figure 8, where the ground station executes the motion planning algorithm, transmits real-time control inputs to the Ackermann UGV, and receives feedback on its configuration from FZMotion motion capture system  [^2] https://ymate3d.com via the Robot Operating System (ROS).
Similar to the setup in the previous section, we prepare two sets of independently conducted examples for the Ackermann UGV. In the first set of three examples, the UGV starts in or passes through the saturation region $\mathcal{S}$, while the UGV remains outside $\mathcal{S}$ for the four examples in the second set. The initial and target configurations are summarized in Table III. To generate control inputs that start from zero, we specify the linear velocity gain by $k_{v}=(1-e^{-0.3t})v_{\text{max}}$, where $t$ is the runtime and other control parameters are consistent with those in the previous section.
@@ VI-A2 Results
The experimental results are presented in Figure 9. Figures 9(a) and LABEL:sub@fig:Exp_inputs4567 show the tracking results of the control inputs. The errors between the current and target configurations are plotted in Figure 9(c). The snapshots of Examples 1-3 and Examples 4-7 are presented in Figures 9(d), LABEL:sub@fig:Exp_half_123, LABEL:sub@fig:Exp_full_123 and 9(e), LABEL:sub@fig:Exp_half_4567, LABEL:sub@fig:Exp_full_4567, respectively. Due to measurement noise, we observe that the tracking results of the linear velocity exhibit small fluctuations around the desired value. In the case of angular velocity, the tracking error increases due to parameter uncertainties and the response time of the UGV’s chassis. However, despite the aforementioned real-world disturbances, the Ackermann UGV consistently converges to the target configuration with bounded trajectory curvature in all examples. The experimental results showcase the effectiveness and robustness of the feedback motion planning algorithm in real-world applications.
(a) Control inputs tracking of Exp 1-3.
@@
Figure 9: Experimental results with the Ackermann UGV. The tracking errors of the control inputs are presented by the shaded areas around the desired values in subfigures LABEL:sub@fig:Exp_inputs_123, LABEL:sub@fig:Exp_inputs4567.
@@ VI-B HIL Experiments with the Fixed-Wing UAV
Figure 10: HIL experiment platform same as [liu2024low] except the planning algorithm.
@@ VI-B1 Setup
To demonstrate that the proposed algorithm, represented by closed-form expressions, can be implemented for real-time onboard computation, we conduct hardware-in-the-loop (HIL) experiments with the fixed-wing UAV. The hardware-in-the-loop (HIL) system is illustrated in Figure 10. In this architecture, the onboard computer manages all the computations of the motion planning algorithm, the autopilot functions as the flight control unit and generates commands for actuators based on our planned control inputs, and the simulator replicates the real-world flight environment while providing configuration feedback.
We consider a fixed-wing UAV with minimum turning radius $\rho=$30\text{\,}\mathrm{m}$\$ and linear velocity bounded by $v_{\text{min}}=$16\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ and $v_{\text{max}}=$18\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Due to the cruising speed constraint, this UAV has positive lower bound on its linear velocity, and hence can only converge to the target positive limit set rather than the target configuration. The parameters of the CVF are specified by $r_{1}=6\rho$, $r_{2}=12\rho$ and $r_{3}=18\rho$.
The HIL experiments are conducted following the procedure outlined below. Initially, the fixed-wing UAV takes off and maneuvers to the initial configuration $\bm{g}_{0}$, at which point the motion planning algorithm is activated, directing the UAV to the target configuration $\bm{g}_{d}$. Once the UAV reaches $\bm{g}_{d}$, the motion planning algorithm is deactivated while the autopilot guides the UAV to the subsequent initial configuration without landing. This process is repeated for all the examples in Table IV, where all the target configurations are encompassed by a common target positive limit set $\mathcal{L}_{d}^{+}=\{\bm{g}\mid r=r_{2},\ \theta=\theta_{r}\}$.
@@ VI-B2 Setpoint Derivation
The control inputs of linear velocity and angular velocity in 2D provided by our algorithm cannot be directly applied to 3D fixed-wing UAVs. Therefore, we first convert the linear and angular velocities into the corresponding setpoints of attitude and thrust for the autopilot platform as follows.
For simplicity, we assume that the attack angle is sufficiently small and that there is no slide-slip angle. Therefore, the body-fixed frame and the velocity frame of the fixed-wing UAV are equivalent. The setpoint for the UAV’s attitude is represented by the yaw-pitch-roll Euler angles, denoted by ($\alpha$, $\beta$, $\gamma$). The attitude matrix $\bm{R}\in\mathrm{SO}(3)$ is given by
@@
where $K_{\beta}$ is the gain, $h$ is the current flight height and $V$ is the airspeed. When the experiment begins, the UAV takes off and then circles in the horizontal plane where $h=h_{d}$, waiting for the motion planning algorithm to take over. In this context, we assume that the pitch angle $\beta$ is already stabilized at zero in the subsequent derivation.
@@
Subsequently, we aim to track the desired angular velocity with the time derivative of the yaw angle. In accordance with the coordinated turn constraint defined by
$\dot{\alpha}=-\frac{a_{g}}{V}\tan{\gamma}\cos{\beta}\overset{\mathrm{\beta=0}}{=}-\frac{a_{g}}{V}\tan{\gamma},$
we derive the setpoint of roll angle that enables $\dot{\alpha}$ to follow $\omega$ as
$\gamma_{d}=-\arctan\frac{\omega V}{a_{g}},$ (29)
where $a_{g}$ is the gravitational acceleration and $\omega$ is determined by (22). Finally, to track the planned linear velocity $v_{x}$ given by (17) with airspeed $V$, we assign the thrust by